1.McKinley Company owns 15,000 of the 50,000 outstanding shares of Ranier Corporation common stock but cannot exert significant influence over Ranier. During 2009, Ranier earns $350,000 and pays cash dividends of $140,000.

a. Indicate the effect on McKinley's 2009 Net Income?

b.what is the net change in McKinley's Investment account during 2009?

2.McKinley Company owns 15,000 of the 50,000 outstanding shares of Ranier Corporation common stock but can exert significant influence over Ranier. During 2009, Ranier earns $350,000 and pays cash dividends of $140,000.

a. Indicate the effect on McKinley's 2009 Net Income?

b.what is the net change in McKinley's Investment account during 2009?
Answer #1

1.

Because McKinley Company cannot exert significant influence over Ranier, it will report the dividend received from Ranier in its income statement as dividend revenue. Ranier's earnings would have no effect on McKinley's net income.

Therefore,

a. McKinley's 2009 Net Income would be affected by $140,000 x 15/50 = 42,000

b. There would be no change in McKinley's Investment account during 2009. The answer would be $0.

2.

Because McKinley Company can exert significant influence over Ranier, it will report the investment in Ranier using equity method. Under this method, McKinley Company will report it share in Ranier's earnings as investment revenue on the income statement.

Therefore,

a. McKinley's 2009 Net Income would be affected by $350,000 x 15/50 = 105,000

b. Net change in McKinley's Investment account during 2009 = 105,000 - 42,000 = 63,000